OpenStudy (anonymous):

In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10 a. x + 2 b.x + 6 c.–x + 2 d. x + 2

OpenStudy (anonymous):

Substitution method?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is confusing to me because it says to solve for y for x – 3y = –6 If you do this you get x – 3y = –6 x +6 = 3y \[y=\frac{x}{3}+2\]

OpenStudy (anonymous):

im sorry these are the answers 2x + y = 11 5x – 2y = –1 a. 2x + 11 b.–2x + 11 c. 2x – 11 d.–2x – 11

OpenStudy (anonymous):

2x + y = 11 2x -2x + y = 11 -2x